Graph the following including the axis of symmetry, identify the axis of symmetry, and create a suitable table of values x on one side and y on the other.
y = 2x^2

Answer by stanbon(48558)

Graph the following including the axis of symmetry, identify the axis of symmetry, and create a suitable table of values x on one side and y on the other.
y = 2x^2
If x=0, y=1
If x=1 or x=-1, y=2
If x=2 or x=-2, y=8
If x=3 or x=-3, y=18
--------------------
The vertex is at x=-b/2a=-0/2*2=0
So the axis of symmetry is x=0
--------------
Cheers,
Stan H.